OpenStudy (anonymous):

simplify. Assume that all variables are positive. 3√54y10

OpenStudy (adamaero):

what number squared equals 54?

OpenStudy (anonymous):

I'm not sure .-.

OpenStudy (anonymous):

It would be a really long decimal

OpenStudy (adamaero):

that's for Nat to answer

OpenStudy (anonymous):

well technically there is but, it is an very long decimal (that is my correction

OpenStudy (adamaero):

haven't you heard of the Socratic method?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

it is not a square root, it is a cubed root

OpenStudy (anonymous):

gonna let admaero handle this. Luck mate.

OpenStudy (anonymous):

\[\sqrt[3]{54y^{10}}\]

OpenStudy (adamaero):

Socrates the person even?

OpenStudy (anonymous):

\[\sqrt[3]{54y^{10}}=\sqrt[3]{27\times 2\times y^9\times y}\] is a start

OpenStudy (anonymous):

nope, my teacher doesn't really teach, she gives us homework and expects us to understand it

OpenStudy (anonymous):

then break it up as \[\sqrt[3]{27}\sqrt[3]{y^9}\sqrt[3]{2y}\]

OpenStudy (anonymous):

thanks satellite

OpenStudy (adamaero):

ah, well it would be a similar question either way I don't know about you satellite73; you know your stuff, but give too close to the answer too soon

OpenStudy (anonymous):

the first two you know you know that \[\sqrt[3]{27}=3\] because \(3^3=27\)

OpenStudy (anonymous):

and you also know that \[\sqrt[3]{y^9}=y^3\]because \[(y^3)^3=y^9\]

OpenStudy (anonymous):

so your final answer is to put it all together get \[\large 3y^3\sqrt[3]{2y}\]

OpenStudy (anonymous):

hope the method was clear (socratic or not...)

OpenStudy (anonymous):

yes it was. thank you again, i appreciate it

OpenStudy (perl):

the socratic method is a bit vague to me

OpenStudy (perl):

if you mean , good guiding question type learning , ok